38.)If the last worker employed ($10/hour) produces 50 additional units of output per hour and the last unit of capital employed ($500/hour) produces 2,500 units of output per hour, how can the firm minimize the cost of its current output?

a.)hire more labor and less capital
b.)hire more capital and less labor
c.)hire more labor and capital
d.)hire less labor and capital
e.)maintain current employment levels of labor and capital


39.)The least-cost hiring rule states that costs are minimized when
a.)each input has an equivalent MP.
b.)a firm employs workers to the point where MRP = MRC.
c.)MP per dollar is equal for all inputs.
d.)VMP is maximized.
e.)MR = MC.
Answer #1

38) e

MPl/w= 50/10= 5

MPk/r= 2500/500= 5

Since both ratios are equal cost is minimized

39) c

At cost minimization MP per dollar is equal for all inputs